the line segment joining the points (-5 -3) and (3 4) is to be divided into five equal parts. Find the point of division closest to (-5 -3)
Found 2 solutions by stanbon, Fombitz:
Answer by stanbon(75887)

he line segment joining the points (-5 -3) and (3 4) is to be divided into five equal parts. Find the point of division closest to (-5 -3)
-------
Let that point be (x,y)
x = -5+(1/5)(3--5) = -5 + (1/5)(8) = -5+(1 3/5) = -3 2/5 = -17/5
y = -3+(1/5)(4--3) = -3+7/5 = -3+(1 2/5) = -1 3/5 = -8/5
-------------
Cheers,
Stan H.
----------------

Answer by Fombitz(32388)